A train slows from 108 km per hour with a uniform acceleration of 5 m per second square. How long will it take to reach 18 km per hour and the what is the distance covered.

1
Expert's answer
2021-10-31T18:12:57-0400

By defintion, the acceleration is given as follows:


"a = \dfrac{v_f - v_i}{t}"

where "v_f = 18km\/h = 5m\/s, v_i = 108km\/h = 30m\/s" are the final and initial speeds respectively, and "t" is the time required to reach from "v_i" to "v_f". Solving for "t" and substituting "a = -5m\/s^2" (minus since it decelerates), find:


"t = \dfrac{v_f-v_i}{a} = \dfrac{30m/s - 5m/s}{-5m/s^2} =5s"

The distance is given by the kinematic equation:


"d = v_it + \dfrac{at^2}{2} = 30m/s\cdot 5s - \dfrac{5m/s^2\cdot (5s)^2}{2} = 87.5m"

Answer. 5s and 87.5m.